OpenStudy (anonymous):

how to factor 2x^2-5x-3=0

rvc (rvc):

\[2x^2-5x-3=0\]

OpenStudy (cj49):

\[-b \pm \sqrt{b ^{2}-4ac}/2\]

OpenStudy (cj49):

a=2 b=-5 c=-3 Plug in the values

rvc (rvc):

also u can use 2 and 3

OpenStudy (anonymous):

my question is how to go from 2x^2-5x-3=0 to (2x+1)(x-3)=0

OpenStudy (cj49):

\[-5+\sqrt{-5^{2}-4(2)(-3)}/2(2)\]

OpenStudy (cj49):

\[-5+\sqrt{25+24}/4\]

rvc (rvc):